OpenStudy (anonymous):

simplify (4x^2-6x+1)+(-2x^2-5x)

OpenStudy (anonymous):

you add like terms

OpenStudy (anonymous):

2x^2-11x+1

OpenStudy (anonymous):

4x^2-2x^2 = 2x^2 -6x-5x=-11x 1+0=1